what is the present value annuity of 1500 payeble at the end of each 6 month period for 2 years if money with 8% compounded semi-annually.
Answer by jorel1380(3719)

The typical annuity payment formula is thus: P= r(PV)/1-(1+r)^-n, where P=payment, PV=present value, r= interest rate per period, and n= total number of periods. Solving for present value, we get:
P[(1-(1+r)^-n)/r]=PV
So:
1500[1-(1+.04)^-4/.04]=PV
PV=5445 is your present value. ☺☺☺☺